Before Heating: Mass of 250 mL beaker and oil 76.552 grams
Mass of beaker + oil +20 kernels 79.691 grams
After Heating: Mass of beaker + oil + popped kernels 79.102 grams
Using the data above, calculate the initial mass of the popcorn kernels and the final mass of the popcorn kernels.

Answer by josgarithmetic(39617)

Mass of the kernels: 9.691-6.552=3.139 grams.

Mass of the loss: 9.691-9.102=0.508 grams.

Final mass of the kernels: 3.139-0.508=2.631 grams; but note that this may be ignoring any absorption of oil into the kernels.